### v2

# A Theoretical Framework for Data Efficient Multi-Source Transfer Learning Based on Cramér-Rao Bound

Abstract
Multi-source transfer learning provides an effective solution to data scarcity in real-world supervised learning scenarios by leveraging multiple source tasks. In this field, existing works typically use all available samples from sources in training, which constrains their training efficiency and may lead to suboptimal results. To address this, we propose a theoretical framework that answers the question: what is the optimal quantity of source samples needed from each source task to jointly train the target model? Specifically, we introduce a generalization error measure that aligns with cross-entropy loss, and minimize it based on the Cramér-Rao Bound to determine the optimal transfer quantity for each source task. Additionally, we develop an architecture-agnostic and data-efficient algorithm OTQMS to implement our theoretical results for training deep multi-source transfer learning models. Experimental studies on diverse architectures and two real-world benchmark datasets show that our proposed algorithm significantly outperforms state-of-the-art approaches in both accuracy and data efficiency. The code and supplementary materials are available in https://anonymous.4open.science/r/Materials.
Machine Learning, ICML
1 Introduction
Nowadays, various machine learning algorithms have achieved remarkable success by leveraging large-scale labeled training data. However, in many practical scenarios, the limited availability of labeled data presents a significant challenge, where transfer learning emerges as an effective solution (Zhao et al., 2020). Transfer learning aims to leverage knowledge from tasks with abundant data or being well-trained, known as the source tasks, to improve the performance of a new learning task, known as the target task. Given its numerous applications, transfer learning has gained wide popularity and seen success in a variety of fields, such as computer vision (Wang & Deng, 2018), natural language processing (Sung et al., 2022), recommendation system (Fu et al., 2024) and anomaly detection (Vincent et al., 2020). Traditionally, transfer learning has focused on the transfer between a single source task and a target task. However, there is a growing emphasis on multi-source transfer learning, which leverages multiple source tasks to enhance the training of the target task (Sun et al., 2015).
Figure 1: More source samples does not always mean better performance. Incorporating all source samples may bring negative impact, which is illustrated by the comparison of two strategies, using target with all source samples (blue) and using target only (red), evaluated on the equally divided 5-task CIFAR10 dataset. Theoretically, although incorporating more source samples reduces model variance by expanding the training data, the discrepancy between the source and target tasks introduces additional bias.
Table 1: Comparison across matching-based transfer learning method, based on whether they are tailored to multi-sources, have task generality, have shot generality, and require target labels. The ‘✓’ represents obtaining the corresponding aspects, while ‘✗’ the opposite. Task generality denotes the ability to handle various target task types, and shot generality denotes the ability to avoid negative transfer in different target sample quantity settings including few-shot and non-few-shot.
Method Multi-Source Task Generality Shot Generality Target Label
MCW (Lee et al., 2019) Supervised
Leep (Nguyen et al., 2020) Supervised
Tong (Tong et al., 2021) Supervised
DATE (Han et al., 2023) Unsupervised
H-ensemble (Wu et al., 2024) Supervised
OTQMS (Ours) Supervised
In multi-source transfer learning, traditional methods usually jointly train the target model using all available samples from sources without selection (Zhang et al., 2024; Shui et al., 2021; Li et al., 2021b). This evidently poses a severe limitation to training efficiency, considering the vast number of available samples from various potential source tasks in real-world scenarios (Peng et al., 2019). Moreover, directly assuming the use of all available samples seriously constrains their solution space, which possibly leads to suboptimal results as illustrated in Figure 1. Therefore, it is critical to establish a theoretical framework to answer the question: what is the optimal transfer quantity of source samples from each source task needed in training the target task?
In this work, we establish a theoretical framework for the multi-source transfer learning problem. We formulate the transfer quantity of each source task as an optimization variable and propose a method to determine its optimal value. Specifically, we introduce the expectation of Kullback-Leibler (K-L) divergence between the distribution learned from training samples and the true distribution of target task samples as a measure of generalization error. This measure is then minimized based on the Cramér-Rao Bound to derive the optimal transfer quantity of each source task. Building on this, we propose a practical algorithm, named Optimal Transfer Quantities for Multi-Source Transfer learning (OTQMS), to implement our theoretical results for training deep multi-source transfer learning models. Notably, OTQMS is data-efficient and compatible with various model architectures, including Vision Transformer (ViT) and Low-Rank Adaptation (LoRA). It also demonstrates advantages in task generality and shot generality (as illustrated in Table 1), since we establish our theoretical framework without restricting it to a specific target task type or limiting the range of target sample quantity.
In summary, our main contributions are as follows:
  • a)
    A novel theoretical framework that optimizes transfer quantities is introduced for multi-source transfer learning. We propose a new K-L divergence measure of the generalization error. By minimizing it, we develop a method to determine the optimal transfer quantity of each source task.
  • b)
    Based on the framework, we propose OTQMS, an architecture-agnostic and data-efficient algorithm for target model training in multi-source transfer learning.
  • c)
    Experimental studies on two real-world datasets and various model architectures demonstrate that OTQMS achieves significant improvements in both accuracy and data efficiency. In terms of accuracy, OTQMS outperforms state-of-the-art by an average of $1.5\%$ on DomainNet and $1.0\%$ on Office-Home. In terms of data efficiency, OTQMS reduces the average training time by $35.19\%$ and the average sample usage by $47.85\%$ on DomainNet. Furthermore, extensive supplementary experiments demonstrate that our theory can facilitate both full model training and parameter-efficient training.
2 Related Work
2.1 Transfer Learning Theory
Existing theoretical works can be categorized into two groups. The first group focuses on proposing measures to quantify the similarity between the target and source tasks. Within this group, some measures have been introduced, including $l_{2}$-distance (Long et al., 2014), optimal transport cost (Courty et al., 2016), LEEP (Nguyen et al., 2020), Wasserstein distance (Shui et al., 2021), and maximal correlations (Lee et al., 2019). This work belongs to the second group focusing on developing new generalization error measures. Within this group, the measures having been introduced include $f$-divergence (Harremoës & Vajda, 2011), mutual information  (Bu et al., 2020), ${\mathcal{X}}^{2}$-divergence (Tong et al., 2021), $\mathcal{H}$-score (Bao et al., 2019; Wu et al., 2024). However, the potential of K-L divergence as a generalization error measure has not been sufficiently explored.
2.2 Multi-source Transfer Learning
Classified by the object of transfer, existing multi-source transfer learning methods mainly focus on two types: model transfer vs sample transfer (Zhuang et al., 2020). Model transfer assumes there is one or more pre-trained models on the source tasks and transfers their parameters to the target task via fine-tuning (Wan et al., 2022). This work focuses on the latter, which is based on joint training of the source task samples with those of the target task (Zhang et al., 2024; Shui et al., 2021; Li et al., 2021b). Classified by the strategy of transfer, existing methods mainly focus on two types: alignment strategy and matching-based strategy (Zhao et al., 2024). Alignment strategy aims to reduce the domain shift among source and target domains (Li et al., 2021a; Zhao et al., 2021; Li et al., 2021b). This work is more similar to the latter, focusing on determining which source domains or samples should be selected or assigned higher weights for transfer (Guo et al., 2020; Shui et al., 2021; Tong et al., 2021; Wu et al., 2024). However, most existing works either utilize all samples from all sources or perform task-level selection, whereas this work explores a framework that optimizes the transfer quantity of each source task.
3 Problem Formulation
Consider the transfer learning setting with one target task ${\mathcal{T}}$, and $K$ source tasks $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$. The target task $\mathcal{T}$ is not restricted to a specific downstream task category. Generally, we formulate it as a parameter estimation problem under a distribution model $P_{X;{\underline{\theta}}}$. For example, when ${\mathcal{T}}$ is a supervised classification task, $P_{X;{\underline{\theta}}}$ corresponds to the joint distribution model of input features $Z$ and output labels $Y$, i.e., $X=(Z,Y)$. Our objective is to estimate the true value of $\underline{\theta}$, which corresponds to optimizing the neural network parameters for target task ${\mathcal{T}}$. Here, $\theta$ denotes 1-dimensional parameter, and $\underline{\theta}$ denotes high dimensional parameter.
Furthermore, we assume that the source tasks and the target task follow the same parametric model and share the same input space $\mathcal{X}$. Without loss of generality, we assume $\mathcal{X}$ to be discrete, though our results can be readily extended to continuous spaces. The target task ${\mathcal{T}}$ has $N_{0}$ training samples $X^{N_{0}}=\{x_{1},\dots,x_{N_{0}}\}$ i.i.d. generated from some underlying joint distribution $P_{X;{\underline{\theta}}_{0}}$, where the parameter ${\underline{\theta}}_{0}\in\mathbb{R}^{d}$. Similarly, the source task $\mathcal{S}_{k}$ has $N_{k}$ training samples $X^{N_{k}}=\{x^{k}_{1},\dots,x^{k}_{N_{k}}\}$ i.i.d. generated from some underlying joint distribution $P_{X;{\underline{\theta}}_{k}}$, where $k\in[1,K]$, and the parameter ${\underline{\theta}}_{k}\in\mathbb{R}^{d}$. In this work, we use the Maximum Likelihood Estimator (MLE) to estimate the true target task parameter $\theta_{0}$. Moreover, the following lemma demonstrates that, for large sample size, the MLE can achieve the theoretical lower bound of the Mean Squared Error (MSE) asymptotically, known as the Cramér-Rao Bound.
Lemma 3.1.
(Cover & Thomas, 2006) Let $\hat{{\underline{\theta}}}$ be an unbiased estimator of ${\underline{\theta}}_{0}$ based on $n$ i.i.d. samples. Then, the MSE matrix of $\hat{{\underline{\theta}}}$ satisfies the following lower bound:
$\text{MSE}(\hat{{\underline{\theta}}})^{d\times d}\triangleq\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-{\underline{\theta}}_{0}\right)\left(\hat{{\underline{\theta}}}-{\underline{\theta}}_{0}\right)^{T}\right]\geq\frac{1}{n}J({\underline{\theta}}_{0})^{-1},$ (1)
where $\left(\hat{{\underline{\theta}}}-{{\underline{\theta}}_{0}}\right)$ is $d-dimensional$ vector, and the matrix inequality $A\geq B$ means that $A-B$ is positive semi-definite. In addition, the $J({\underline{\theta}})$ is fisher information matrix defined as:
$J({\underline{\theta}})^{d\times d}$ $=\mathbb{E}\left[\left(\frac{\partial}{\partial{\underline{\theta}}}\log P(X;{\underline{\theta}})\right)\left(\frac{\partial}{\partial{\underline{\theta}}}\log P(X;{\underline{\theta}})\right)^{T}\right].$ (2)
Moreover, as the sample size $n\to\infty$, the MLE $\hat{{\underline{\theta}}}$ can achieve this bound asymptotically, i.e.,
$\text{MSE}(\hat{{\underline{\theta}}})=\frac{1}{n}J({\underline{\theta}}_{0})^{-1}+o\left(\frac{1}{n}\right).$ (3)
It should be noted that in some works, the MSE is defined as a scalar, which corresponds to the trace of the MSE matrix and can also be bounded using this lemma.
When we transfer $n_{1},\dots,n_{K}$ samples from the $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$, where $n_{k}\in\left[0,N_{k}\right]$, we denote these training sample sequences as $X^{n_{1}},X^{n_{2}},\dots,X^{n_{k}}$ Then, the MLE for multi-source transfer learning, in the simplest form, is a summation of the likelihood of all source tasks and target tasks, i.e.,
$\hat{{\underline{\theta}}}=\operatorname*{arg\,max}_{{\underline{\theta}}}$ ${\sum_{x\in X^{N_{0}}}}\log P_{X;{\underline{\theta}}}(x)$
$+{\sum_{k=1}^{K}\sum\limits_{x\in X^{n_{k}}}}\log P_{X;{\underline{\theta}}}(x).$ (4)
In this work, our goal is to derive the optimal transfer quantities $n_{1}^{*},\dots,n_{K}^{*}$ of source tasks $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$ to minimize certain divergence measure $\mathcal{L}$ between the distribution learned from training samples $P_{X;\hat{{\underline{\theta}}}}$ and the true distribution of target task $P_{X;{\underline{\theta}}_{0}}$, i.e.,
$n_{1}^{*},\dots,n_{k}^{*}=\operatorname*{arg\,min}_{n_{1},\dots,n_{k}}\mathcal{L}(P_{X;\hat{{\underline{\theta}}}},P_{X;{\underline{\theta}}_{0}}).$ (5)
Besides, we provide the notations table in Appendix A.
4 Theoretical Analysis and Algorithm
In this section, we will first introduce a new K-L divergence measure for the optimization problem in (5). Then, we will analyze it based on the Cramér-Rao bound to derive the optimal transfer quantities for both single-source and multi-source scenarios. Finally, we will develop a practical algorithm based on the theoretical framework.
Definition 4.1.
(The K-L divergence) The K-L divergence measures the difference between two probability distributions $P(X)$ and $Q(X)$ over the same probability space. It is defined as:
$D\left(P\middle\|Q\right)=\sum_{x\in\mathcal{X}}p(x)\log\frac{p(x)}{q(x)}.$
In this work, we apply the K-L divergence between the distribution model parameterized by MLE (3) and the true distribution model of target task to measure the generalization error, which directly corresponds to the testing error using cross-entropy loss (Cover & Thomas, 2006).
$D\left(P_{X;\hat{{\underline{\theta}}}}\middle\|P_{X;{\underline{\theta}}_{0}}\right)=-{\underbrace{H(P_{X;\hat{{\underline{\theta}}}})}_{\textnormal{entropy}}}+\underbrace{\mathbb{E}_{P_{X;\hat{{\underline{\theta}}}}}\left[P_{X;{\underline{\theta}}_{0}}\right]}_{\textnormal{cross-entropy}}.$
Finally, our generalization error measure is defined as follows.
$\mathcal{L}(P_{X;\hat{{\underline{\theta}}}},P_{X;{\underline{\theta}}_{0}})=\mathbb{E}\left[D\left(P_{X;\hat{{\underline{\theta}}}}\middle\|P_{X;{\underline{\theta}}_{0}}\right)\right].$ (6)
4.1 Single-Source Transfer Learning
The direct computation of the proposed K-L divergence measure is challenging. Fortunately, we can prove that the proposed K-L measure directly depends on the MSE in the asymptotic case. Moreover, the MSE has a widely used Cramér-Rao bound introduced in Lemma 3.1, which provides a bound on the estimator’s error, and is asymptotically tight. As our goal is to find a correspondence between transfer quantity and generalization error, for the rest of the paper, we mainly analyze the K-L measure based on the Cramér-Rao bound (3). To begin, we consider the setting with a target task ${\mathcal{T}}$ with $N_{0}$ samples generated from a model with 1-dimensional parameter.
Lemma 4.2.
(proved in Appendix B.1) Given a target task ${\mathcal{T}}$ with $N_{0}$ i.i.d. samples generated from 1-dimensional underlying model $P_{X;\theta_{0}}$, where $\theta_{0}\in\mathbb{R}$, and denoting $\hat{\theta}$ as the MLE (3) based on the $N_{0}$ samples, then the K-L measure (6) can be expressed as:
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}}\middle\|P_{X;\theta_{0}}\right)\right]=\frac{1}{2N_{0}}+o\left(\frac{1}{N_{0}}\right).$ (7)
The result of Lemma 4.2 demonstrates that when there is only a target task, our K-L measure is inversely proportional to the number of training samples. Next, we consider the transfer learning scenario in which we have one target task ${\mathcal{T}}$ with $N_{0}$ training samples and one source task $\mathcal{S}_{1}$ with $N_{1}$ training samples. In this case, We aim to determine the optimal transfer quantity $n_{1}^{*}\in[1,N_{1}]$. To facilitate our mathematical derivations, we assume $N_{0}$ and $N_{1}$ are asymptotically comparable, and the distance between the parameters of target task and source task is sufficiently small (i.e., $|\theta_{0}-\theta_{1}|=O(\frac{1}{\sqrt{N_{0}}})$). Considering the similarity of low-level features among tasks of the same type, this assumption is made without loss of generality. Furthermore, as demonstrated in subsequent analysis, our conclusions remain valid even in extreme cases where the distance between parameters is large.
Theorem 4.3.
(proved in Appendix B.2) Given a target task ${\mathcal{T}}$ with $N_{0}$ i.i.d. samples generated from 1-dimensional underlying model $P_{X;\theta_{0}}$, and a source task $\mathcal{S}_{1}$ with $N_{1}$ i.i.d. samples generated from 1-dimensional underlying model $P_{X;\theta_{1}}$, where $\theta_{0},\theta_{1}\in\mathbb{R}$. $\hat{\theta}$ is denoted as the MLE (3) based on the $N_{0}$ samples from ${\mathcal{T}}$ and $n_{1}$ samples from $\mathcal{S}_{1}$, where $n_{1}\in[0,N_{1}]$, then the K-L measure (6) can be expressed as:
[…]
where
$t\triangleq J(\theta_{0})(\theta_{1}-\theta_{0})^{2}.$ (9)
Moreover, the optimal transfer quantity $n_{1}^{*}$ is
$n_{1}^{*}=\begin{cases}N_{1},&\quad\mathrm{if}\ N_{0}\cdot t\leq 0.5\\ \min\left(N_{1},\frac{N_{0}}{2N_{0}t-1}\right),&\quad\mathrm{if}\ N_{0}\cdot t>0.5\end{cases}.$ (10)
From (8), we observe that the K-L measure decreases as $N_{0}$ increases, which aligns with our intuition that utilizing all available target samples is beneficial. In addition, the trend of (8) with respect to $n_{1}$ is more complex. We plot (8) as a function of $n_{1}$ under two different regimes, determined by the value of $N_{0}\cdot t$, as shown in Figure 2. Our goal is to explore the optimal value of $n_{1}^{*}$ to minimize (8).
Figure 2: The function curve figures of (8) under different regimes determined by the value of $N_{0}\cdot t$ (blue). The vertical axis denotes the value of K-L measure (8), while the horizontal axis denotes the variable $n_{1}$.
  • Case 1 ($N_{0}\cdot t\leq 0.5$): The K-L measure monotonically decreases as $n_{1}$ increases. Obviously, the optimal point is $n_{1}^{*}=N_{1}$. This indicates that when the source task and the target task are highly similar, i.e., when $t$ is small, an increase in the transfer quantity will positively impact the results.
  • Case 2 ($N_{0}\cdot t>0.5$): The K-L measure first decreases and then increases as $n_{1}$ increases. It attaining its minimum at $n_{1}^{\prime}=\frac{N_{0}}{2N_{0}t-1}$. It should be noted that when $t$ is large enough, the point $n_{1}^{\prime}$ approaches 00. This aligns with the intuition that when the discrepancy between the source and target tasks is substantial, avoiding transfer yields better results. Furthermore, if $n_{1}^{\prime}$ exceeds $N_{1}$, we should utilize all $N_{1}$ samples, so $n_{1}^{*}=\min\left(N_{1},\frac{N_{0}}{2N_{0}t-1}\right)$.
In Figure 3, we provide an intuitive explanation for the optimal $n_{1}^{*}$.
Figure 3: The three circles represent the errors corresponding to transfer quantity $n_{1}=0$, $n_{1}^{*}$, and $N_{1}$. The distance from each circle’s center to $\theta_{0}$ represents the bias of estimation, and the radius represents the variance. As the transfer quantity increases, the bias increases, while the variance decreases. The optimal $n_{1}^{*}$ achieves the best trade-off between them.
As the dimensionality of the model parameter $\theta$ increases to higher dimensions, we derive the following propositions.
Proposition 4.4.
(proved in Appendix B.3) In the case where the parameter dimension is d, i.e., ${\underline{\theta}}_{0}\in\mathbb{R}^{d}$, with all other settings remaining the same as the Lemma 4.2, the K-L measure (6) can be expressed as
$\frac{d}{2N_{0}}+o\left(\frac{1}{N_{0}}\right).$ (11)
Proposition 4.5.
(proved in Appendix B.4) In the case where the parameter dimension is d, i.e., ${\underline{\theta}}_{0},{\underline{\theta}}_{1}\in\mathbb{R}^{d}$, with all other settings remaining the same as the Theorem 4.3, the K-L measure (6) can be expressed as
$\frac{d}{2}\left(\frac{1}{N_{0}+n_{1}}+\frac{n_{1}^{2}}{(N_{0}+n_{1})^{2}}t\right)+o\left(\frac{1}{N_{0}+n_{1}}\right),$ (12)
where we denote
$t\triangleq\frac{({{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}})^{T}J({{\underline{\theta}}_{0}})({{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}})}{d}.$ (13)
In addition, t is a scalar, ${J}({{\underline{\theta}}_{0}})$ is $d\times d$ matrix, and $\left({{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}}\right)$ is a $d-dimensional$ vector, which is the element-wise subtraction of two d-dimensional vectors ${\underline{\theta}}_{1}$ and ${\underline{\theta}}_{0}$.
Compared to Theorem 4.3, Proposition 4.5 exhibits a similar mathematical form, allowing us to derive the optimal transfer quantity through a similar approach. Furthermore, we observe that as the parameter dimension $d$ increases, the K-L error measure (12) increases. This suggests that for a complex model, knowledge transfer across tasks becomes more challenging, which is consistent with the findings in (Tong et al., 2021).
Table 2: Multi-Source Transfer Performance on DomainNet and Office-Home. The arrows indicate transfering from the rest tasks. The highest/second-highest accuracy is marked in Bold/Underscore form respectively.
Method Backbone DomainNet Office-Home
[…]
Unsupervised-all-shots
MSFDA(Shen et al., 2023) ResNet50 66.5 21.6 56.7 20.4 70.5 54.4 48.4 75.6 62.8 84.8 85.3 77.1
DATE(Han et al., 2023) ResNet50 - - - - - - - 75.2 60.9 85.2 84.0 76.3
M3SDA(Peng et al., 2019) ResNet101 57.2 24.2 51.6 5.2 61.6 49.6 41.5 - - - - -
Supervised-10-shots
Few-Shot Methods:
H-ensemble(Wu et al., 2024) ViT-S 53.4 21.3 54.4 19.0 70.4 44.0 43.8 71.8 47.5 77.6 79.1 69.0
MADA(Zhang et al., 2024) ViT-S 51.0 12.8 60.3 15.0 81.4 22.7 40.5 78.4 58.3 82.3 85.2 76.1
MADA(Zhang et al., 2024) ResNet50 66.1 23.9 60.4 31.9 75.4 52.5 51.7 72.2 64.4 82.9 81.9 75.4
MCW(Lee et al., 2019) ViT-S 54.9 21.0 53.6 20.4 70.8 42.4 43.9 68.9 48.0 77.4 86.0 70.1
WADN(Shui et al., 2021) ViT-S 68.0 29.7 59.1 16.8 74.2 55.1 50.5 60.3 39.7 66.2 68.7 58.7
Source-Ablation Methods:
Target-Only ViT-S 14.2 3.3 23.2 7.2 41.4 10.6 16.7 40.0 33.3 54.9 52.6 45.2
Single-Source-Avg ViT-S 50.4 22.1 44.9 24.7 58.8 42.5 40.6 65.2 53.3 74.4 72.7 66.4
Single-Source-Best ViT-S 60.2 28.0 55.4 28.4 66.0 49.7 48.0 72.9 60.9 80.7 74.8 72.3
AllSources $\cup$ Target ViT-S 71.7 32.4 60.0 31.4 71.7 58.5 54.3 77.0 62.3 84.9 84.5 77.2
OTQMS (Ours) ViT-S 72.8 33.8 61.2 33.8 73.2 59.8 55.8 78.1 64.5 85.2 84.9 78.2
4.2 Multi-Source Transfer Learning
Consider the multi-source transfer learning scenario with $K$ source task $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$ and one target task ${\mathcal{T}}$. We aim to derive the optimal transfer quantity $n_{i}^{*}$ of each source.
Theorem 4.6.
(proved in Appendix B.5) Given a target task ${\mathcal{T}}$ with $N_{0}$ i.i.d. samples generated from underlying model $P_{X;{\underline{\theta}}_{0}}$, and K source tasks $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$ with $N_{1},\dots,N_{K}$ i.i.d. samples generated from underlying model $P_{X;{\underline{\theta}}_{1}},\dots,P_{X;{\underline{\theta}}_{K}}$, where ${\underline{\theta}}_{0},{\underline{\theta}}_{1},\dots,{\underline{\theta}}_{K}\in\mathbb{R}^{d}$. $\hat{{\underline{\theta}}}$ is denoted as the MLE (3) based on the $N_{0}$ samples from ${\mathcal{T}}$ and $n_{1},\dots,n_{K}$ samples from $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$, where $n_{i}\in[0,N_{i}]$. Denoting $s=\sum\limits_{i=1}^{K}n_{i}$ as the total transfer quantity, and $\alpha_{i}=\frac{n_{i}}{s}$ as the proportion of different source tasks, then the K-L measure (6) can be expressed as:
$\frac{d}{2}\left(\frac{1}{N_{0}+s}+\frac{s^{2}}{(N_{0}+s)^{2}}t\right)+o\left(\frac{1}{N_{0}+s}\right).$ (14)
In (14), $t$ is a scalar denoted as
$t=\frac{\underline{\alpha}^{T}\Theta^{T}{J}({{\underline{\theta}}_{0}})\Theta\underline{\alpha}}{d},$ (15)
where $\underline{\alpha}=\left[\alpha_{1},\dots,\alpha_{K}\right]^{T}$ is a K-dimensional vector, and $\Theta^{d\times K}=\left[{{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}},\dots,{{\underline{\theta}}_{K}}-{{\underline{\theta}}_{0}}\right]$.
According to Theorem 4.6, we can derive the optimal transfer quantities $n_{1}^{*},\dots,n_{K}^{*}$ by minimizing (14). Equivalently, we need to find the optimal total transfer quantity $s^{*}$ and the optimal proportion vector $\underline{\alpha}^{*}$ which minimize (14). The analytical solutions of $s^{*}$ and $\underline{\alpha}^{*}$ are difficult to acquire, and we provide a method to get their numerical solutions in Appendix D. Eventually, we can get the optimal transfer quantity of each source through $n_{i}^{*}=s^{*}\cdot\alpha_{i}^{*}$
4.3 Practical Algorithm
Along with our theoretical framework, we propose a practical algorithm, OTQMS, which is applicable to all supervised target tasks, as presented in Algorithm 1. It mainly involves two stages: (1) initializing the source task parameters $\theta_{1},...,\theta_{K}$ to compute the optimal transfer quantities and (2) jointly training the target model using a resampled dataset whose sample quantity of each source corresponds to the optimal transfer quantity derived from our theory.
Algorithm 1 OTQMS: Training
1:  Input: Target data $D_{{\mathcal{T}}}=\{(z_{\mathcal{T}}^{i},y_{\mathcal{T}}^{i})\}_{i=1}^{N_{0}}$, source data $\{D_{S_{k}}=\{(z_{S_{k}}^{i},y_{S_{k}}^{i})\}_{i=1}^{N_{k}}\}_{k=1}^{K}$, model type $f_{{\underline{\theta}}}$ and its parameters ${\underline{\theta}}_{0}$ for target task and $\{{\underline{\theta}}_{k}\}_{k=1}^{K}$ for source tasks, parameter dimension $d$. // $z$ represents the feature and $y$ represents the label
2:  Parameter: Learning rate $\eta$.
3:  Initialize: randomly initialize ${\underline{\theta}}_{0}$, $\{{\underline{\theta}}_{k}\}_{k=1}^{K}$.
4:  Output: a well-trained model $f_{{\underline{\theta}}_{0}}$ for target task.
5:  for $k=1$ to $K$ do // Initialize the source parameters
6:      repeat
7:         $L_{k}\leftarrow\frac{1}{N_{k}}\sum\limits_{i=1}^{N_{k}}\ell\left(y_{S_{k}}^{i},f_{{\underline{\theta}}_{k}}(z_{S_{k}}^{i})\right)$
8:         ${\underline{\theta}}_{k}\leftarrow{\underline{\theta}}_{k}-\eta\nabla_{{\underline{\theta}}_{k}}L_{k}$
9:      until ${\underline{\theta}}_{k}$ converges;
10:  end for $D_{train}\leftarrow D_{\mathcal{T}}=\{(z^{i},y^{i})\}_{i=1}^{N_{train}}$
11:  repeat // Joint training of the target task
12:     $L_{{\mathcal{T}}}\leftarrow\frac{1}{N_{train}}\sum\limits_{i=1}^{N_{train}}\ell\left(y^{i}_{{\mathcal{T}}},f_{{\underline{\theta}}_{0}}(z^{i})\right)$
13:     ${\underline{\theta}}_{0}\leftarrow{\underline{\theta}}_{0}-\eta\nabla_{{\underline{\theta}}_{0}}L_{{\mathcal{T}}}$
14:     $\Theta\leftarrow\left[{{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}},\dots,{{\underline{\theta}}_{K}}-{{\underline{\theta}}_{0}}\right]^{T}$
15:     ${J}({{\underline{\theta}}_{0}})\leftarrow(\nabla_{{\underline{\theta}}_{0}}L_{{\mathcal{T}}})(\nabla_{{\underline{\theta}}_{0}}L_{{\mathcal{T}}})^{T}$
16:     $(s^{*},\underline{\alpha}^{*})\leftarrow\operatorname*{arg\,min}\limits_{(s,\underline{\alpha})}\frac{d}{2}\left(\frac{1}{N_{0}+s}+\frac{s^{2}}{(N_{0}+s)^{2}}\frac{\underline{\alpha}^{T}\Theta^{T}{J}({{\underline{\theta}}_{0}})\Theta\underline{\alpha}}{d}\right)$
17:     $D_{source}\leftarrow\bigcup\limits_{k=1}^{K}\left\{D_{S_{k}}^{*}\bigg{|}D_{S_{k}}^{*}\subseteq D_{S_{k}},\;\lvert D_{S_{k}}^{*}\rvert=s^{*}\alpha_{k}^{*}\right\}$
18:     $D_{train}\leftarrow D_{source}\bigcup D_{\mathcal{T}}$
19:  until ${\underline{\theta}}_{0}$ converges;
5 Experiments
5.1 Experiments Settings
Benchmark Datasets. DomainNet contains 586,575 samples of 345 classes from 6 domains (i.e., C: Clipart, I: Infograph, P: Painting, Q: Quickdraw, R: Real and S: Sketch). Office-Home benchmark contains 15588 samples of 65 classes, with 12 adaptation scenarios constructed from 4 domains: Art, Clipart, Product and Real World (abbr. Ar, Cl, Pr and Rw). Digits contains four-digit sub-datasets: MNIST(mt), Synthetic(sy), SVHN(sv) and USPS(up), with each sub-dataset containing samples of numbers ranging from 0 to 9.
Implementation Details. We employ the ViT-Small model (Wightman, 2019), pre-trained on ImageNet-21k (Deng et al., 2009), as the backbone for all datasets. The Adam optimizer is employed with a learning rate of $1e^{-5}$. We allocate 20% of the dataset as the test set, and report the highest accuracies within 5 epoch early stops in all experiments. Following the standard few-shot learning protocol, the training data for k-shot consists of k randomly selected samples per class from the target task. All experiments are conducted on Nvidia A800 GPUs.
Baselines. For a general performance evaluation, we take SOTA works under similar settings as baselines. The scope of compared methods includes: 1) Unsupervised Methods: MSFDA (Shen et al., 2023), DATE (Han et al., 2023), M3SDA (Peng et al., 2019). 2) Few-Shot Methods Based on Model(Parameter)-Weighting: H-ensemble (Wu et al., 2024), MCW (Lee et al., 2019). 3) Few-Shot Methods Based on Sample: MADA (Zhang et al., 2024), WADN (Shui et al., 2021) 4) Source Ablating Methods: Target-Only, Single-Source-Avg/Single-Source-Best (average/best performance of single-source transfer), AllSources $\cup$ Target (all source & target data in multi-source transfer). Note that MADA (Zhang et al., 2024) leverages all unlabeled target data in conjunction with a limited amount of labeled target data, which is a hybrid approach combining unsupervised and supervised learning. Due to the page limit, we provide detailed information on the experimental design and the results of an experiment adapted to the WADN settings on the Digits dataset in Appendix C.2.
5.2 Main Result
We evaluated our algorithm, OTQMS, alongside baseline methods on the multi-source transfer learning tasks using the DomainNet and Office-Home datasets. The quantitative results are summarized in Table 2. Since the unsupervised baselines are not designed for the supervised few-shot setting, we report their original results from the respective papers for reference. We make the following observations:
Overall Performance. In general, compared to baseline methods, OTQMS achieves the best performance on almost all the transfer scenarios on the two datasets. Specifically, OTQMS outperforms the state-of-the-art (AllSources $\cup$ Target) by an average of $1.5\%$ on DomainNet and $1.0\%$ on Office-Home.
Data Speaks Louder Than Model Weighting. It is worth noting that on both datasets, samples-based methods utilizing both target and source samples to jointly train the model, such as WADN, MADA and OTQMS, generally outperform model(parameter)-weighting approaches which construct the target model by weighting source models, such as H-ensemble and MCW. This observation suggests that sample-based approaches offer greater advantages over model-based methods, because they can fully leverage the relevant information from the source data for the target task.
Take, But Only as Much as You Need. Comparing results in Table 2 among Target-only, AllSources $\cup$ Target, and OTQMS, we observe that OTQMS achieves the best performance in both datasets by leveraging only a subset of data selected from all available sources based on model preference. This result validates our theory and highlights the benefits of identifying the optimal balance in the bias-variance trade-off. By choosing the right quantities of samples from the source tasks, we could train the target model more accurately. Furthermore, Figure 5 shows that OTQMS also significantly reduces the training time and sample usage, which validates its superiority in terms of data efficiency.
Few-Shot Labels, Big Gains. We make a comparison of the results of unsupervised and supervised methods. While other conditions remain the same, Table 2 demonstrates that even if unsupervised methods like MSFDA and M3SDA take all the target data into account (up to 1.3×$10^{5}$ samples on Real domain of DomainNet), their performance still falls short compared to the supervised methods, which rely on only a limited number of samples (3450 samples). This illustrates the importance of having supervised information in multi-source transfer learning.
5.3 Static vs. Dynamic Transfer Quantity
In our proposed Algorithm 1, we employ a “Dynamic” strategy that dynamically determines the optimal transfer quantities and updates the resampled dataset during the joint training of target task. To validate the effectiveness of this strategy, we conducted comparative experiments using the “Static-*” methods. “Static-*” methods first simulate the distribution of target on target dataset only, and different types of Static such as “Under, Exact and Over” stands for different fitting levels. In “Static-*” methods, we only compute the optimal transfer quantity once to make the resampled dataset, and evaluated on it until target model converges. The results on Table 3 demonstrate OTQMS using dynamic transfer quantity achieved the best performance.
Table 3: Static vs. Dynamic Transfer Quantity in OTQMS on Office-Home. Arrows will transfer from the rest tasks.
Method Backbone Office-Home
[…]
Supervised-10-shots:
Static-Under ViT-S 77.0 62.3 84.9 84.5 77.2
Static-Exact ViT-S 46.0 59.8 85.1 83.7 68.7
Static-Over ViT-S 76.8 61.9 78.6 68.6 71.5
Dynamic ViT-S 78.1 64.5 85.2 84.9 78.2
5.4 Generality across Different Shot Settings
As discussed in the theoretical analysis of Theorem 4.3, our theoretical framework is applicable to any quantity of target samples. Therefore, OTQMS exhibits shot generality, enabling it to avoid negative transfer across different shot settings. To validate this, we increase the number of shots from 5 to 100 across methods including AllSources $\cup$ Target, Target-Only, and OTQMS. As shown in Figure 4, experimental results demonstrate that OTQMS consistently outperforms other approaches across all shot settings. This highlights the generality and scalability of OTQMS in terms of data utilization.
Figure 4: Performance comparison with increasing target shots up to 100 per class on DomainNet dataset (I, P, Q and R domains). OTQMS (blue) outperforms other methods.
5.5 Data Efficiency Test
In this section, we demonstrate the advantage of OTQMS in terms of data efficiency. Specifically, we conduct experiments with MADA, AllSources $\cup$ Target, and OTQMS across different shot settings, and for each shot setting, we accumulate the total sample used and time consumed until the highest accuracy is reached. To better visualize the results, we present the average sample usage and training time across all shot settings in Figure 5. To be specific, OTQMS reduces the average training time by $35.19\%$ and the average sample usage by $47.85\%$ on DomainNet, compared to the AllSources $\cup$ Target method.
Figure 5: Data efficiency comparison of average sample usage and training time on DomainNet dataset, the left vertical axis represents the amount of sample usage, with green bars indicating AllSources $\cup$ Target data counts, blue bars about OTQMS, red bars about MADA(ViT-S) and azury bars about MADA(Res50), while the right orange vertical axis and lines represent training time.
5.6 Compatibility to Parameter-Efficient Training
To evaluate the applicability with parameter-efficient training, we used a ViT-Base model (Dosovitskiy et al., 2021) integrated with the LoRA framework (Hu et al., 2021) as the backbone. This approach significantly reduces the number of trainable parameters for downstream tasks while ensuring model quality and improving training efficiency. In the experiments, we consider the trainable low-rank matrices and the classification head as the parametric model in our theoretical framework, while treating the remaining parameters as constants. Other experimental settings are the same as default. Experiments were conducted on the Office-Home dataset. The results of Table 5 in Appendix demonstrate that our method remains effective.
5.7 Analysis on Domain-specific Transfer Quantity
To understand the domain preference of OTQMS, we visualize the proportion of each source in the selected samples for each target domain. As shown in Figure 6(b), when the target domain is Clipart, OTQMS primarily leverages samples from Real, Painting, and Sketch. In addition, Quickdraw contributes minimally to any target domain. These observations align with the findings in (Peng et al., 2019). To further clarify the selection process of OTQMS during training, we visualize it in Figure 6(a). We observe that in the Office-Home dataset, Clipart initially selects all source domains but later focuses on Art and Real World. In the DomainNet dataset, Sketch predominantly selects Clipart, Painting, and Real throughout the training process.
Refer to captionRefer to caption
(a) Domain Selection
(b) Domain preference
Figure 6: Visualization of domain-specific transfer quantity under 10-shot setting. (a) Domain selection during training epochs (from left to right), where the blue upper part represents the selection of target domain Clipart on Office-Home, and the orange lower part represents the selection of target domain Sketch on DomainNet. Darker colors indicate stronger tendencies throughout the training process. (b) Source domain preferences of different target domains on DomainNet. Each row corresponds to a target domain while each column represents a source domain.
6 Conclusion
In this work, we propose a theoretical framework to determine the optimal transfer quantities in multi-source transfer learning. Our framework reveals that by optimizing the transfer quantity of each source task, we can improve target task training while reducing the total transfer quantity. Based on this theoretical framework, we develop an architecture-agnostic and data-efficient practical algorithm OTQMS for jointly training the target model. We evaluated the proposed algorithm through extensive experiments and demonstrated its superior accuracy and enhanced data efficiency.
lmpact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Notations
      Symbol       Description       ${\mathcal{T}}$       target task       $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$       source tasks       $N_{0}$        quantity of target samples       $N_{1}\cdots N_{k}$       maximum sample quantity of each source       $n_{1}\cdots n_{k}$       transfer quantity of each source       $\theta$       model parameter       $\underline{\theta}$       vectorized model parameter       $\underline{\theta}_{0}$       vectorized model parameter of target task       $\underline{\theta}_{i},i\in[1,K]$       vectorized model parameter of i-th source task,       $J(\theta)$       Fisher information (scalar) of $\theta$       $J(\underline{\theta})^{d\times d}$       Fisher information (matrix) of d-dimensional $\underline{\theta}$       $P(X;\theta)$       probability of X parameterized by $\theta$       $P_{X;\theta}$       distribution of X parameterized by $\theta$       $||\underline{x}||^{2}$       l-2 norm of vector x       $\alpha_{i}$       transfer proportion in multi-source case from i-th source task       $\underline{\alpha}$       transfer proportion vector in multi-source case whose i-th entry is $\alpha_{i}$       $s$       total transfer quantity in multi-source case       $\hat{\theta}$       estimator of $\theta$       $E_{\hat{\theta}}$       expectation of $\hat{\theta}$
Table 4: Notations
Appendix B Proofs
B.1 Proof of Lemma 4.2
Lemma B.1.
In the asymptotic case, the K-L measure (6) and the mean squared error have the relation as follows.
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}}\middle\|P_{X;\theta_{0}}\right)\right]$
$=\frac{1}{2}{J}(\theta_{0})\text{MSE}(\hat{\theta})+o(\frac{1}{N_{0}}),$ (16)
Proof.
In this section, for the sake of clarity, we will write $\hat{\theta}$ in its parameterized form $\hat{\theta}(X^{N_{0}})$ when necessary, and these two forms are mathematically equivalent. By taking Taylor expansion of $D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)$ at $\theta_{0}$, we can get
$D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)=\frac{1}{2}\sum_{x\in X}\frac{\left(P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})$ (17)
We denote $\delta$ as a small constant, and we can rewrite (6) as
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)\right]$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(P_{X;\hat{\theta}(X^{N_{0}})}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})\right)$
$=\sum_{X^{N_{0}}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})\right)$
=$=\sum_{\left\{X^{N_{0}}:\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}<\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})\right)$
+$+\sum_{\left\{X^{N_{0}}:\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}\geq\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})\right)$
To facilitate the subsequent proof, we introduce the concept of ”Dot Equal”.
Definition B.2.
(Dot Equal ($\dot{=}$)) Specifically, given two functions $f(n)$ and $g(n)$, the notation $f(n)\dot{=}g(n)$ is defined as
$f(n)\dot{=}g(n)\quad\Leftrightarrow\quad\lim_{n\to\infty}\frac{1}{n}\log\frac{f(n)}{g(n)}=0,$ (19)
which shows that $f(n)$ and $g(n)$ have the same exponential decaying rate.
We denote $\hat{P}_{X^{N_{0}}}$ as the empirical distribution of $X^{N_{0}}$. Applying Sanov’s Theorem to (B.1), we can know that
$P_{X^{n};\theta_{0}}(X^{N_{0}})\doteq e^{-N_{0}D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)}$ (20)
Then, we aim to establish a connection between the K-L divergence (20) and $\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}$. From (17), we can know that
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)=\frac{1}{2}\sum_{x\in X}\frac{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})$ (21)
From the features of MLE, we can know that
$\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\hat{\theta}(X^{N_{0}})}}(x)}{\partial\hat{\theta}}\right]=0=\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}\right]+\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]{\left(\hat{\theta}(X^{N_{0}})-\theta_{0}\right)},$ (22)
which can be transform to
$\left(\hat{\theta}(X^{N_{0}})-\theta_{0}\right)=\frac{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}\right]}{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]}=\frac{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\frac{\partial P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}}{P_{X;{\theta_{0}}}(x)}\right]}{\mathbb{E}_{\hat{P}_{X^{N_{0}}}}\left[\frac{\partial^{2}\log P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}^{2}}\right]}=\frac{\sum\limits_{x\in\mathcal{X}}{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)}\frac{\frac{\partial P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}}{P_{X;{\theta_{0}}}(x)}}{J(\theta_{0})}$ (23)
Using the Cauchy-Schwarz inequality, we can obtain
$\sum_{x\in X}\frac{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)^{2}}{P_{X;\theta_{0}}(x)}\cdot\sum_{x}\frac{(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}})^{2}}{P_{X;\theta_{0}}(x)}\geq\left(\sum\limits_{x\in\mathcal{X}}{\left(\hat{P}_{X^{N_{0}}}(x)-P_{X;\theta_{0}}(x)\right)}\frac{\frac{\partial P_{X;{\theta_{0}}}(x)}{\partial\theta_{0}}}{P_{X;{\theta_{0}}}(x)}\right)^{2}$ (24)
where
$\sum_{x}\frac{(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}})^{2}}{P_{X;\theta_{0}}(x)}=\sum_{x}P_{X;\theta_{0}}(x)\left(\frac{1}{P_{X;\theta_{0}}(x)}\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}\right)^{2}=\sum_{x}P_{X;\theta_{0}}(x)\left(\frac{\partial\log P_{X;\theta_{0}}(x)}{\partial\theta_{0}}\right)^{2}=J(\theta_{0})$ (25)
Combining with (21), (23), and (25), the inequality (24) can be transformed to
$D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)\geq\frac{J(\theta_{0})\left(\hat{\theta}(X^{N_{0}})-\theta_{0}\right)^{2}}{2}=\frac{J(\theta_{0})\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}}{2}$ (26)
Combining (20) and (26), we can know that
$P_{X^{n};\theta_{0}}(X^{N_{0}})\doteq e^{-N_{0}D\left(\hat{P}_{X^{N_{0}}}\middle\|P_{X;\theta_{0}}\right)}\leq e^{\frac{-N_{0}J(\theta_{0})\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}}{2}}$ (27)
For the first term in (B.1), the $\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|$ is small enough for us to omit the term $o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})$. As for the second term in in (B.1), even though the magnitude of $\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|$ is no longer negligible, the probability of such sequences is $O(e^{\frac{-N_{0}J(\theta_{0})\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}}{2}})$ by (27), which is exponentially decaying with $N_{0}$ such that the second term is $o{(\frac{1}{N_{0}})}$. By transfering (B.1), we can get
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}(X^{N_{0}})}\middle\|P_{X;\theta_{0}}\right)\right]$
=$=\sum_{\left\{X^{N_{0}}:\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2}<\delta\right\}}P_{X^{n};\theta_{0}}(X^{N_{0}})\left(\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right)+o(\|\hat{\theta}(X^{N_{0}})-\theta_{0}\|^{2})$
[…]
$=\frac{1}{2}\mathbb{E}\left[\frac{1}{2}\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right]+o{(\frac{1}{N_{0}})}.$ (28)
We then transform (B.1) with (25)
$\frac{1}{2}\mathbb{E}\left[\sum_{x\in X}\frac{\left(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}}(\hat{\theta}(X^{N_{0}})-\theta_{0})\right)^{2}}{P_{X;\theta_{0}}(x)}\right]$
$=\frac{1}{2}\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]\sum_{x}\frac{(\frac{\partial P_{X;\theta_{0}}(x)}{\partial\theta_{0}})^{2}}{P_{X;\theta_{0}}(x)}$
$=\frac{1}{2}\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]{J}(\theta_{0})$ (29)
Combining (B.1), (B.1), we can get
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}}\middle\|P_{X;\theta_{0}}\right)\right]=\frac{1}{2}\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]{J}(\theta_{0})+o{(\frac{1}{N_{0}})}.$ (30)
From (B.1), we can establish the relationship between the proposed K-L measure (6) and the mean squared error. Then, by using the (3), the K-L measure is
$\frac{1}{2N_{0}}+o{(\frac{1}{N_{0}})}.$ (31)
B.2 Proof of Theorem 4.3
In this section we will be using samples from both distributions for our maximum likelihood estimation, so our optimization problem becomes
$\hat{\theta}=\operatorname*{arg\,max}_{\theta}L_{n}(\theta),$ (32)
where, when using $N_{0}$ samples from target task, and $n_{1}$ samples from source task
$L_{n}(\theta)\triangleq\frac{1}{N_{0}+n_{1}}\sum_{x\in X^{N_{0}}}\log P_{X;\theta}(x)+\frac{1}{N_{0}+n_{1}}\sum_{x\in X^{n_{1}}}\log P_{X;\theta}(x).$ (33)
And, we also define the expectation of our estimator, which is somewhere between $\theta_{0}$ and $\theta_{1}$ to be $E_{\hat{\theta}}$
$E_{\hat{\theta}}=\operatorname*{arg\,max}_{\theta}L(\theta),$ (34)
We could equivalently transform (34) into
$E_{\hat{\theta}}=\operatorname*{arg\,min}_{\theta}\frac{N_{0}D\left(P_{X;\theta_{0}}\middle\|P_{X;\theta}\right)}{N_{0}+n_{1}}+\frac{n_{1}D\left(P_{X;\theta_{1}}\middle\|P_{X;\theta}\right)}{N_{0}+n_{1}}$ (35)
By taking argmin of (35) we can get
$P_{X;E_{\hat{\theta}}}=\frac{N_{0}P_{X;\theta_{0}}+n_{1}P_{X;\theta_{1}}}{N_{0}+n_{1}},$ (36)
By doing a Taylor Expansion of (36) around $\theta^{\prime}$, which is in the neighbourhood of $\theta_{1},\theta_{2}$ and $E_{\hat{\theta}}$, we can get
$E_{\hat{\theta}}=\frac{N_{0}\theta_{0}+n_{1}\theta_{1}}{N_{0}+n_{1}}+O\left(\frac{1}{N_{0}+n_{1}}\right),$ (37)
Lemma B.3.
[…]
Proof.
Since $\hat{\theta}$ is a maximizer of $L_{n}{(\theta)}$, $L_{n}^{{}^{\prime}}{(\hat{\theta})}=0=L_{n}^{{}^{\prime}}{(E_{\hat{\theta}})}+L_{n}^{{}^{\prime\prime}}{(E_{\hat{\theta}})}(\hat{\theta}-E_{\hat{\theta}})$. Therefore,
$\sqrt{N_{0}+n_{1}}(\hat{\theta}-E_{\hat{\theta}})=\frac{\sqrt{N_{0}+n_{1}}L_{n}^{{}^{\prime}}{(E_{\hat{\theta}})}}{L_{n}^{{}^{\prime\prime}}{(E_{\hat{\theta}})}}.$ (39)
Since $E_{\hat{\theta}}$ maximizes $L{(\theta)}$, $L^{{}^{\prime}}{(E_{\hat{\theta}})}=\frac{N_{0}}{N_{0}+n_{1}}\mathbb{E}_{\theta_{0}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial x}\right]+\frac{n_{1}}{N_{0}+n_{1}}\mathbb{E}_{\theta_{1}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x_{i}^{{}^{\prime}})}}{\partial x}\right]=0$. Therefore,
$\sqrt{N_{0}+n_{1}}L_{n}^{{}^{\prime}}(E_{\hat{\theta}})$
$=\sqrt{\frac{N_{0}}{N_{0}+n_{1}}}\left(\sqrt{\frac{1}{N_{0}}}\sum_{x\in X^{N_{0}}}\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}-\sqrt{N_{0}}\mathbb{E}_{\theta_{0}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right]\right)$
$+\sqrt{\frac{n_{1}}{N_{0}+n_{1}}}\left(\sqrt{\frac{1}{n_{1}}}\sum_{x\in X^{n_{1}}}\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}-\sqrt{n_{1}}\mathbb{E}_{\theta_{1}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right]\right)$ (40)
Applying the Central Limit Theorem to (B.2), we can get
$\sqrt{N_{0}+n_{1}}L_{n}^{{}^{\prime}}(E_{\hat{\theta}})\xrightarrow{a.s.}\mathcal{N}\Bigg{(}0,$ $\frac{N_{0}}{N_{0}+n_{1}}\left(\mathbb{E}_{\theta_{0}}\left[\left(\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)^{2}\right]-\mathbb{E}_{\theta_{0}}\left[\left(\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)\right]^{2}\right)+$
$\frac{n_{1}}{N_{0}+n_{1}}\left(\mathbb{E}_{\theta_{1}}\left[\left(\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)^{2}\right]-\mathbb{E}_{\theta_{1}}\left[\left(\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)\right]^{2}\right)\Bigg{)}$ (41)
By taking Taylor expansion of $E_{\hat{\theta}}$ at $\theta_{1}$, we can get
$\mathbb{E}_{\theta_{0}}\left[\left(\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)^{2}\right]$
$=\mathbb{E}_{\theta_{0}}\left[\left(\frac{\partial\log{P_{X;\theta_{0}}(x)}}{\partial\theta}+\frac{\partial}{\partial\theta}\frac{\partial\log{P_{X;\theta_{0}}(x)}}{\partial\theta}(E_{\hat{\theta}}-\theta_{0})+O(\frac{1}{N_{0}+n_{1}})\right)^{2}\right]$
$=J(\theta_{0})+(E_{\hat{\theta}}-\theta_{0})\mathbb{E}_{\theta_{0}}\left[\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\frac{\partial^{2}\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta^{2}}\right]+O(\frac{1}{N_{0}+n_{1}})$
$=J(\theta_{0})+O(\frac{1}{\sqrt{N_{0}+n_{1}}})$ (42)
and
$\mathbb{E}_{\theta_{0}}\left[\left(\frac{\partial\log{P_{X;E_{\hat{\theta}}}(x)}}{\partial\theta}\right)\right]^{2}$
$=\mathbb{E}_{\theta_{0}}\left[\left(\frac{\partial\log{P_{X;\theta_{0}}(x)}}{\partial\theta}+\frac{\partial}{\partial\theta}\frac{\partial\log{P_{X;\theta_{0}}(x)}}{\partial\theta}(E_{\hat{\theta}}-\theta_{0})+O(\frac{1}{N_{0}+n_{1}})\right)\right]^{2}$
$=\mathbb{E}_{\theta_{0}}\left[\left(\frac{\partial}{\partial\theta}\frac{\partial\log{P_{X;\theta_{0}}(x)}}{\partial\theta}(E_{\hat{\theta}}-\theta_{0})+O(\frac{1}{N_{0}+n_{1}})\right)\right]^{2}$
$=(E_{\hat{\theta}}-\theta_{0})^{2}E_{\theta_{0}}\left[\left(\frac{\partial}{\partial\theta}\frac{\partial\log{P_{X;\theta_{0}}(x)}}{\partial\theta}\right)\right]^{2}+o(\frac{1}{N_{0}+n_{1}})$
$=O(\frac{1}{N_{0}+n_{1}})$ (43)
By combining (B.2), (B.2), and (B.2), we can get
$\operatorname{var}\left(\sqrt{N_{0}+n_{1}}L_{n}^{{}^{\prime}}(E_{\hat{\theta}})\right)=\frac{N_{0}}{N_{0}+n_{1}}J(\theta_{0})+\frac{n_{1}}{N_{0}+n_{1}}J(\theta_{1})+O(\frac{1}{\sqrt{N_{0}+n_{1}}})$ (44)
[…]
Under the assumption that $\theta_{0},\theta_{1},E_{\hat{\theta}}$ are sufficiently close to each other, we can easily deduce that the difference among ${J}(\theta_{0}),{J}(\theta_{1})$ and ${J}(E_{\hat{\theta}})$ is $O(\frac{1}{\sqrt{N_{0}+n_{1}}})$. We can easily get
[…]
Therefore, the bound of $\mathbb{E}\left[\left(\hat{\theta}-E_{\hat{\theta}}\right)^{2}\right]$ is
$\frac{1}{(N_{0}+n_{1})J(\theta_{0})}$ (47)
Combining (6) and (B.1), we know that
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}}\middle\|P_{X;\theta_{0}}\right)\right]$
$=\frac{1}{2}{J}(\theta_{0})\mathbb{E}\left[\left(\hat{\theta}-\theta_{0}\right)^{2}\right]+o(\frac{1}{N_{0}})$
$=\frac{1}{2}{J}(\theta_{0})\bigg{(}\mathbb{E}\left[\left(\hat{\theta}-E_{\hat{\theta}}\right)^{2}\right]+\mathbb{E}\left[2\left(\hat{\theta}-E_{\hat{\theta}}\right)\left(E_{\hat{\theta}}-\theta_{0}\right)\right]+\mathbb{E}\left[\left(E_{\hat{\theta}}-\theta_{0}\right)^{2}\right]\bigg{)}+o(\frac{1}{N_{0}})$
$=\frac{1}{2}{J}(\theta_{0})\left(\mathbb{E}\left[\left(\hat{\theta}-E_{\hat{\theta}}\right)^{2}\right]+\mathbb{E}\left[\left(E_{\hat{\theta}}-\theta_{0}\right)^{2}\right]\right)+o(\frac{1}{N_{0}})$ (48)
Combining (B.2), (37) and (47), we know that the K-L measure is
$\frac{1}{2}\frac{1}{N_{0}+n_{1}}+\frac{1}{2}{J}(\theta_{0})\frac{n_{1}^{2}}{(N_{0}+n_{1})^{2}}\left(\theta_{1}-\theta_{0}\right)^{2}$ (49)
B.3 Proof of Proposition 4.4
Similar to (B.1), we can get
$\mathbb{E}\left[D\left(P_{X;\hat{{\underline{\theta}}}}\middle\|P_{X;{\underline{\theta}}_{0}}\right)\right]$
$=\frac{1}{2}tr\left({J}({\underline{\theta}}_{0})\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-{\underline{\theta}}_{0}\right)\left(\hat{{\underline{\theta}}}-{\underline{\theta}}_{0}\right)^{T}\right]\right)+o(\frac{1}{N_{0}})$
Combining with Lemma  3.1, we will know that the K-L measure is
$\frac{1}{2}tr\left({J}({\underline{\theta}}_{0})\left({J}({\underline{\theta}}_{0})^{-1}\frac{1}{N_{0}}+o(\frac{1}{N_{0}})\right)\right)=\frac{d}{2N_{0}}+o(\frac{1}{N_{0}})$ (51)
B.4 Proof of Proposition 4.5
Similar to (B.2), we can get
$\mathbb{E}\left[D\left(P_{X;\hat{{\underline{\theta}}}}\middle\|P_{X;{\underline{\theta}}_{0}}\right)\right]$
$=\frac{1}{2}tr\left({J}({\underline{\theta}}_{0})\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-{\underline{\theta}}_{0}\right)\left(\hat{{\underline{\theta}}}-{\underline{\theta}}_{0}\right)^{T}\right]\right)+o(\frac{1}{N_{0}})$
$=\frac{1}{2}\left(tr\left({J}({\underline{\theta}}_{0})\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)^{T}\right]\right)+tr\left({J}({\underline{\theta}}_{0})\mathbb{E}\left[\left(E_{\hat{{\underline{\theta}}}}-{\underline{\theta}}_{0}\right)\left(E_{\hat{{\underline{\theta}}}}-{\underline{\theta}}_{0}\right)^{T}\right]\right)\right)+o(\frac{1}{N_{0}})$ (52)
Similar to (38),we can get
[…]
So we can know that $\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)^{2}\right]$ has the bound
$\frac{1}{(N_{0}+n_{1})}J({\underline{\theta}}_{0})^{-1}$ (54)
The same to (37), we can get
$E_{\hat{{\underline{\theta}}}}=\frac{N_{0}{\underline{\theta}}_{0}+n_{1}{\underline{\theta}}_{1}}{N_{0}+n_{1}}+O\left(\frac{1}{N_{0}+n_{1}}\right),$ (55)
Combining (B.4), (54) and (55), we can know that the K-L measure is
$\frac{d}{2}\left(\frac{1}{N_{0}+n_{1}}+\frac{n_{1}^{2}}{(N_{0}+n_{1})^{2}}t\right)+o\left(\frac{1}{N_{0}+n_{1}}\right),$ (56)
where we denote
$t\triangleq\frac{({{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}})^{T}J({{\underline{\theta}}_{0}})({{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}})}{d}.$ (57)
B.5 Proof of Theorem 4.6
Proof.
Similar to (38),we can get
[…]
So we can know that $\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-E_{\hat{{\underline{\theta}}}}\right)^{2}\right]$ has the bound
$\frac{1}{(N_{0}+s)}J({\underline{\theta}}_{0})^{-1}$ (59)
Similar to (37),we can get
$E_{\hat{{\underline{\theta}}}}=\frac{N_{0}{\underline{\theta}}_{0}+\sum\limits_{i=1}^{k}n_{i}{\underline{\theta}}_{i}}{N_{0}+\sum\limits_{i=1}^{K}n_{i}}=\frac{N_{0}{\underline{\theta}}_{0}+\sum\limits_{i=1}^{k}n_{i}{\underline{\theta}}_{i}}{N_{0}+s}+O\left(\frac{1}{N_{0}+s}\right)$ (60)
Combining (B.2), (59) and(60) , we can know that the K-L measure is
$\frac{1}{2}\frac{1}{N_{0}+s}+\frac{1}{2}\frac{s^{2}}{(N_{0}+s)^{2}}J({\underline{\theta}}_{0})\left(\sum\limits_{i=1}^{k}\alpha_{i}({\underline{\theta}}_{i}-{\underline{\theta}}_{0})\right)^{2}+o\left(\frac{1}{N_{0}+s}\right)$ (61)
For the d-demension parameter, we can know that the K-L measure is $\underline{\theta}_{i}$
$\frac{d}{2}\left(\frac{1}{N_{0}+s}+\frac{s^{2}}{(N_{0}+s)^{2}d}tr\left(\mathbf{J}(\underline{\theta}_{0})\left(\sum\limits_{i=1}^{k}\alpha_{i}(\underline{\theta}-\underline{\theta}_{0})\right)\left(\sum\limits_{i=1}^{k}\alpha_{i}(\underline{\theta}_{i}-\underline{\theta}_{0})\right)^{T}\right)\right)$
$=\frac{d}{2}\left(\frac{1}{N_{0}+s}+\frac{s^{2}}{(N_{0}+s)^{2}}\frac{\underline{\alpha}^{T}\Theta^{T}{J}(\underline{\theta}_{0})\Theta\underline{\alpha}}{d}\right)+o\left(\frac{1}{N_{0}+s}\right)$
where $\underline{\alpha}=\left[\alpha_{1},\dots,\alpha_{K}\right]^{T}$ is a K-dimensional vector, and
${{\Theta}}^{d\times K}=\left[{{\underline{\theta}}_{1}}-{{\underline{\theta}}_{0}},\dots,{{\underline{\theta}}_{K}}-{{\underline{\theta}}_{0}}\right].$ (63)
Appendix C Experiment Details
C.1 Details information of LoRA framework experiments.
Table 5: Multi-Source Transfer with LoRA on Office-Home. We apply LoRA on ViT-B backbone for PEFT.
Method Backbone Office-Home
[…]
Supervised-10-shots Source-Ablation:
Target-Only ViT-B 59.8 42.2 69.5 72.0 60.9
Single-Source-avg ViT-B 72.2 59.9 82.6 81.0 73.9
Single-Source-best ViT-B 74.4 61.8 84.9 81.9 75.8
AllSources $\cup$ Target ViT-B 81.1 66.0 88.0 89.2 81.1
OTQMS (Ours) ViT-B 81.5 68.0 89.2 90.3 82.3
C.2 Experimental Design and Model Adaptation
To ensure consistency in the experimental setup, we first evaluate the performance of different methods on the DomainNet and Office-Home datasets by adapting their settings to align with ours, such as the backbone, dataset, and early stopping criteria. Specifically, for the MADA method, we adjusted the preset of keeping 5% of labeled target samples to 10-shots per class target samples while maintaining other conditions. And in turn, ours is adapted to the WADN settings, equipped with a 3-layer ConvNet and evaluated on Digits dataset.
Table 6: Performance on Digits Dataset. The arrows indicate transferring from the rest tasks. “3Conv” denotes the backbone with 3 convolution layers.
Method Backbone Digits
[…]
Following settings of WADN:
WADN(Shui et al., 2021) 3Conv 88.3 70.6 81.5 90.5 82.7
AllSources $\cup$ Target 3Conv 92.6 67.1 82.5 88.8 82.8
OTQMS (Ours) 3Conv 93.8 67.1 83.3 89.1 83.3
C.3 Single-Source transfer performance details.
Table 7: Single-Source Transfer Performance on DomainNet. The details accuracy information of the “Single-Source-*” lines of Table 2.
Target Domain Backbone Source Domain
Clipart Infograph Painting Quickdraw Real Sketch Avg Best
Clipart ViT-S - 46.5 55.4 30.3 60.2 59.7 50.4 60.2
Infograph ViT-S 25.6 - 25.3 7.3 28.0 24.4 22.1 28.0
Painting ViT-S 49.6 47.3 - 22.4 55.4 49.6 44.9 55.4
Quickdraw ViT-S 26.9 18.1 23.9 - 25.9 28.4 24.7 28.4
Real ViT-S 64.6 62.2 66.0 38.5 - 62.7 58.8 66.0
Sketch ViT-S 49.7 40.9 48.1 26.1 47.8 - 42.5 49.7
Table 8: Single-Source Transfer Performance on Office-Home. The details accuracy information of the “Single-Source-*” lines of Table 2.
Target Domain Backbone Source Domain
Art Clipart Product Real World Avg Best
Art ViT-S - 61.7 61.1 72.9 65.2 72.9
Clipart ViT-S 49.8 - 49.4 60.9 53.3 60.9
Product ViT-S 68.8 73.7 - 80.7 74.4 80.7
Real World ViT-S 70.9 72.3 74.8 - 72.7 74.8
Table 9: Single-Source Transfer Performance on Office-Home of LoRA. The details accuracy information of the “Single-Source-*” lines of Table 5. ViT-B backbone is already frozen and equipped with small trainable LoRA layers.
Target Domain Backbone Source Domain
Art Clipart Product Real World Avg Best
Art ViT-B - 68.4 74.4 73.8 72.2 74.4
Clipart ViT-B 58.3 - 59.6 61.8 59.9 61.8
Product ViT-B 82.0 80.8 - 84.9 82.6 84.9
Real World ViT-B 81.0 80.3 81.9 - 81.0 81.9
C.4 Baselines experiments settings in Table 2.
Since the significant difference from unsupervised methods and supervised methods, the results of MSFDA(Shen et al., 2023), DATE(Han et al., 2023) and M3SDA(Peng et al., 2019) are directly supported by their own article.
On all the experiments of the baselines, we take all the source samples of different domains from trainset into account. And types of the backbone are all pretrained on ImageNet21k(Deng et al., 2009).
As for based on model-parameter-weighting few-shot methods: MCW(Lee et al., 2019) and H-ensemble(Wu et al., 2024), since they have not taken experiments on DomainNet and Office-Home datasets with ViT-Small backbone and 10-shot per class training samples, we take it by changing the backbone and samples condition while maintaining other configurations supported by their own work. And for fairness and efficiency, the well-trained source models from different domains of these methods are directly equipped with the same models trained by the first stage of our OTQMS method respectively.
As for based on samples few-shot methods: We exam the WADN(Shui et al., 2021) method under the condition of limited label on target domain as this setting is similar to ours. And we also change the backbone and samples condition while maintaining other configurations of its report to realize the experiments on DomainNet and Office-Home datasets. It is apparent that the settings of MADA(Zhang et al., 2024) are quite different of ours, leveraging all the labeled source data and all the unlabeled target data with the few-shot labeled ones that are difficult to classify, which means all the target samples have been learned to some extent. So we not only decrease the labeled target samples to 10-shot per class but also the unlabeled target samples. And since MADA is the most SOTA and comparable method, we realize it with our fair 10-shot setting under both ViT-S and ResNet50 backbone on DomainNet and Office-Home datasets while maintaining other configurations of its report.
C.5 Data efficient test details in 5.5 .
Table 10: Data Efficient Test Results on DomainNet. The capital letters represent the target domains.
Method Backbone Data Counts (×106) Training Consumed Time (×104 Second)
C I P Q R S Avg C I P Q R S Avg
MADA(Zhang et al., 2024) ViT-S 8.70 13.39 8.31 4.02 19.76 21.32 12.58 6.40 12.52 6.29 3.22 13.19 18.18 9.97
MADA(Zhang et al., 2024) ResNet50 7.83 8.21 4.57 3.69 4.35 8.78 6.24 7.56 9.58 4.56 4.05 4.21 9.58 6.59
AllSources $\cup$ Target ViT-S 2.17 1.42 1.42 4.20 1.42 1.90 2.09 0.56 0.37 0.39 1.09 0.36 0.49 0.54
OTQMS (Ours) ViT-S 1.18 1.15 0.89 0.97 1.19 1.16 1.09 0.35 0.33 0.28 0.34 0.47 0.36 0.35
Table 11: Data Efficient Log Scaled Test Results on DomainNet. The capital letters represent the target domains.
Method Backbone Data Counts (log scale) Training Consumed Time (log scale)
C I P Q R S Avg C I P Q R S Avg
MADA(Zhang et al., 2024) ViT-S 15.98 16.41 15.93 15.21 16.80 16.87 16.35 11.07 11.74 11.05 10.38 11.79 12.11 11.51
MADA(Zhang et al., 2024) ResNet50 15.87 15.92 15.33 15.12 15.29 15.99 15.65 11.23 11.47 10.73 10.61 10.65 11.47 11.10
AllSources $\cup$ Target ViT-S 14.59 14.17 14.18 15.25 14.17 14.46 14.55 8.63 8.21 8.21 9.30 8.19 8.50 8.59
OTQMS (Ours) ViT-S 13.98 13.95 13.70 13.78 13.99 13.96 13.90 8.17 8.10 7.93 8.12 8.46 8.18 8.17
C.6 Domain choosing analysis details.
To compute the heatmap matrix visualizing domain preference in Figure 6(b), for each source domain, we count the samples selected from it until the target model converges under the 10-shot condition. Since the quantities of available samples varies significantly across domains, we normalize the counts by these quantities. The final domain preference is then determined by computing the importance of these normalized values.
Similarly, to compute the heatmap matrix for domain selection in Figure 6(a), we calculate the importance of the counts of selected samples from different source domains throughout the training epochs.
Appendix D Method to get $s^{*}$ and $\underline{\alpha}^{*}$ which minimize (14) in Theorem 4.6
The minimization problem of K-L measure (14) is
$(s^{*},\underline{\alpha}^{*})\leftarrow\operatorname*{arg\,min}\limits_{(s,\underline{\alpha})}\frac{d}{2}\left(\frac{1}{N_{0}+s}+\frac{s^{2}}{(N_{0}+s)^{2}}\frac{\underline{\alpha}^{T}\Theta^{T}{J}({{\underline{\theta}}_{0}})\Theta\underline{\alpha}}{d}\right).$ (64)
We decompose this problem and explicitly formulate the constraints as follows.
$(s^{*},\underline{\alpha}^{*})\leftarrow\operatorname*{arg\,min}\limits_{s\in[0,\sum\limits_{i=1}^{K}N_{i}]}\frac{d}{2}\left(\frac{1}{N_{0}+s}+\frac{s^{2}}{(N_{0}+s)^{2}d}\operatorname*{arg\,min}\limits_{\underline{\alpha}\in\mathcal{A}(s)}\underline{\alpha}^{T}\Theta^{T}{J}({{\underline{\theta}}_{0}})\Theta\underline{\alpha}\right),$ (65)
where
$\mathcal{A}(s)=\left\{\underline{\alpha}\Bigg{|}\sum_{i=1}^{K}\alpha_{i}=1,s*\alpha_{i}\leq N_{i},\alpha_{i}\geq 0,i=1,\dots,K\right\}.$ (66)
Due to the complex constraints between $s$ and $\alpha$, obtaining an analytical solution to this problem is challenging. Therefore, we propose a numerical approach to get the optimal solution. Specifically, we perform a grid search over the feasible domain of $s$. For each candidate $s^{\prime}$, we compute the optimal $\alpha^{\prime}$ under the constraint $A(s^{\prime})$, which is a $K\times K$ quadratic programming problem with respect to $\alpha$. The K-L measure (14) value is then evaluated using $(s^{\prime},\alpha^{\prime})$. After completing the search over $s$, the optimal values $s^{*}$ and $\alpha^{*}$ correspond to the pair $(s^{\prime},\alpha^{\prime})$ that minimizes the K-L measure.